# FedVSR: Towards Model-Agnostic Federated Learning in Video Super-Resolution

### v1

@@ 5.3 Ablation Study
To further analyze the impact of the DWT loss term and loss-aware adaptive aggregation, we conducted a series of experiments. Using FedAvg as a baseline, we added each component separately to evaluate their individual effects. To assess the impact of the $\mathcal{L}_{\text{HiFr}}$ term, we applied it to FedAvg and observed the changes (F-$\mathcal{L}_{\text{HiFr}}$). Similarly, to evaluate loss-aware adaptive aggregation, we first trained VSR models using a greedy aggregation strategy (F-GLA) that prioritizes low-loss updates in all rounds. We then compared this approach to an adaptive version that incorporates a decay mechanism (F-ALA). As shown in Table 2, the DWT loss term consistently outperforms loss-aware adaptive aggregation in all cases. Additionally, both methods individually surpass FedAvg [31]. However, in some cases, the greedy aggregation strategy resulted in only marginal improvements or even a drop in overall performance especially in ”more heterogeneous” setting based on PSNR and SSIM metrics. These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights.
@@ 7 Further Analysis: Performance Metrics and Visual Comparisons
This section provides additional performance evaluations for our proposed method. Figures 3, 4, and 5 illustrate the PSNR and SSIM trends across different rounds for multiple test sets under various configurations of VRT [25], RVRT [24], and IART [49]. These results further highlight the effectiveness of our approach in enhancing the performance across different models.
Figures 6, 7, and 8 provide qualitative visual comparisons on various test frames. Each figure presents a reference ground truth (GT) frame alongside low-quality (LQ) input and the outputs generated by different Federated Learning strategies, including FedAvg, FedProx, SCAFFOLD, and our method (FedVSR). The reconstructions show that FedVSR produces clearer and more visually faithful outputs, particularly in preserving finer details and reducing artifacts.
Figure 3: PSNR and SSIM values across different rounds for various test sets under different settings for VRT[25].
Figure 4: PSNR and SSIM values across different rounds for various test sets under different settings for RVRT[24].
Figure 5: PSNR and SSIM values across different rounds for various test sets under different settings for IART[49].
Figure 6: Comparison of RVRT output trained with different Federated Learning algorithms under more heterogeneous setting
Figure 7: Comparison of VRT output trained with different Federated Learning algorithms under more heterogeneous setting
Figure 8: Comparison of IART output trained with different Federated Learning algorithms under less heterogeneous setting

### v2

@@ 5.3 Ablation Study
To further analyze the impact of the DWT loss term and loss-aware adaptive aggregation, we conducted a series of experiments. Using FedAvg as a baseline, we added each component separately to evaluate their individual effects. To assess the impact of the $\mathcal{L}_{\text{HiFr}}$ term, we applied it to FedAvg and observed the changes (F-$\mathcal{L}_{\text{HiFr}}$). Similarly, to evaluate loss-aware adaptive aggregation, we first trained VSR models using a greedy aggregation strategy (F-GLA) that prioritizes low-loss updates in all rounds. We then compared this approach to an adaptive version that incorporates a decay mechanism (F-ALA). As shown in Table 2, the DWT loss term consistently outperforms loss-aware adaptive aggregation in all cases. Additionally, both methods individually surpass FedAvg [36]. However, in some cases, the greedy aggregation strategy resulted in only marginal improvements or even a drop in overall performance especially in ”more heterogeneous” setting based on PSNR and SSIM metrics. These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights. Due to space limitations, LPIPS metric results are provided in the supplementary material 7
@@ 6 Conclusion
7 Extended Evaluation of HiFr Loss and Adaptive Aggregation Using LPIPS
To complement the analysis based on PSNR and SSIM discussed earlier in the paper, we present here an additional evaluation using LPIPS, which serves as a perceptual similarity metric. This acts as an extension to the previously discussed metrics and provides a more complete picture of model performance. Due to space limitations, we include this analysis here.
To evaluate the impact of the DWT loss term and the loss-aware adaptive aggregation strategies, we conducted a set of ablation experiments using FedAvg as the baseline. Each component was added independently to isolate its individual effect. Specifically, we assessed the contribution of the $\mathcal{L}_{\text{HiFr}}$ term by applying it to FedAvg (F-$\mathcal{L}_{\text{HiFr}}$), and examined two aggregation variants: a greedy loss-aware approach (F-GLA) that prioritizes updates from clients with lower losses, and an adaptive version (F-ALA) that incorporates a decay factor to regulate client influence over time.
As shown in Table 3 with LPIPS results, the observed trends remain largely consistent across metrics. The DWT loss term consistently provides more significant gains compared to the aggregation strategies, and both methods individually outperform the standard FedAvg baseline. However, similar to PSNR and SSIM, the greedy strategy sometimes leads to only marginal improvements or even minor performance degradation, especially under high heterogeneity. These findings underscore the importance of high-frequency information not only for boosting perceptual quality but also for reliably identifying valuable client updates.
Model Setting Less Heterogeneous More Heterogeneous
REDS Vid4 REDS Vid4
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
VRT F-$\mathcal{L}_{\text{HiFr}}$ 29.80 0.8543 0.1910 24.91 0.7681 0.2399 28.98 0.8322 0.2130 24.70 0.7577 0.2500
F-ALA 29.75 0.8531 0.1912 24.89 0.7681 0.2411 28.90 0.8313 0.2140 24.64 0.7559 0.2520
F-GLA 29.70 0.8523 0.1914 24.86 0.7670 0.2436 28.83 0.8279 0.2198 24.55 0.7500 0.2555
RVRT F-$\mathcal{L}_{\text{HiFr}}$ 31.42 0.8888 0.1506 26.02 0.8178 0.1910 30.24 0.8617 0.1612 25.84 0.8042 0.1978
F-ALA 31.38 0.8880 0.1518 26.00 0.8171 0.1931 30.19 0.8601 0.1728 25.79 0.8041 0.1992
F-GLA 31.31 0.8879 0.1527 25.95 0.8155 0.1948 29.88 0.8576 0.1807 25.60 0.8032 0.2011
IART F-$\mathcal{L}_{\text{HiFr}}$ 30.79 0.8801 0.1588 25.46 0.8001 0.1999 29.64 0.8508 0.1899 25.21 0.7837 0.2231
F-ALA 30.70 0.8789 0.1591 25.44 0.7998 0.2101 29.60 0.8500 0.1903 25.18 0.7831 0.2246
F-GLA 30.57 0.8730 0.1690 25.42 0.7951 0.2119 29.50 0.8482 0.1929 24.99 0.7761 0.2284
Table 3: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR, SSIM, and predicted LPIPS under different heterogeneity settings.
@@ 8 Further Analysis: Performance Metrics and Visual Comparisons
This section provides additional performance evaluations for our proposed method. Figures 3, 5, and 7 illustrate the PSNR, SSIM and LPIPS trends across different rounds for multiple test sets under various configurations of VRT [30], RVRT [29], and IART [54]. These results further highlight the effectiveness of our approach in enhancing the performance across different models.
Figures 4, 6, and 8 provide qualitative visual comparisons on various test frames. Each figure presents a reference ground truth (GT) frame alongside low-quality (LQ) input and the outputs generated by different Federated Learning strategies, including FedAvg, FedProx, SCAFFOLD, and our method (FedVSR). The reconstructions show that FedVSR produces clearer and more visually faithful outputs, particularly in preserving finer details and reducing artifacts.
Figure 3: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for VRT.
Refer to caption (a) Frame 000, Clip 015, REDS
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 32.41, 0.8981, 0.1456 Refer to caption (d) FedProx[27] 32.10, 0.8935, 0.1511
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 28.85, 0.7625, 0.1934 Refer to caption (g) FedVSR (Ours) 33.79, 0.9169, 0.1151
Figure 4: Comparison of RVRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
Figure 5: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for RVRT.
Refer to caption (a) Frame 040, Clip 1, Vid4
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 21.15, 0.7185, 0.3520 Refer to caption (d) FedProx[27] 20.43, 0.6735, 0.4101
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 20.41, 0.6679, 0.3803 Refer to caption (g) FedVSR (Ours) 21.19, 0.7200, 0.3433
Figure 6: Comparison of VRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
Figure 7: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for IART.
Refer to caption (a) Frame 070, Clip 000, REDS
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 27.87, 0.8211, 0.2228 Refer to caption (d) FedProx[27] 27.54, 0.8132, 0.2247
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 27.69, 0.8151, 0.2207 Refer to caption (g) FedVSR (Ours) 28.07, 0.8319, 0.2014
Figure 8: Comparison of IART output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.